OpenStudy (anonymous):

f'(x)= 200 - 1600/(x + 8) - x

OpenStudy (anonymous):

...You want to find f(x), the antiderivative? f(x)=200x-1600ln(x+8)-(1/2)x^2

OpenStudy (anonymous):

no the derivative

OpenStudy (anonymous):

f ' (x)

OpenStudy (anonymous):

You yourself already wrote "f'(x)=200-1600/(x+8)-x", which is the derivative of f(x). The derivative of f'(x) is f''(x), which is f''(x)=1600/(x+8)^2-1.

OpenStudy (anonymous):

i wrote f (x) but want to the deriv f ' (x) of that function

OpenStudy (anonymous):

not the 2nd deriv

OpenStudy (anonymous):

I quote you, "f'(x)= 200 - 1600/(x + 8) - x". That's what you originally wrote. f'(x).

OpenStudy (anonymous):

ah whatever!! lets just find the deriv

OpenStudy (anonymous):

I already gave it, then, except wrote it as f''(x). The derivative of the function you gave me is (d/dx)(200-1600/(x+8)-x)=1600/(x+8)^2-1